OpenStudy (anonymous):

square root{9x+67} = x + 5 I know the solution is 6, however i am stumped trying to figure out how to start. My apologies, i do not know exactly how to write square root signs, so whatever is in those funny looking parenthesis is what's inside the square root sign.

TheSmartOne (thesmartone):

\(\sf\Large\sqrt{9x+67} = x + 5\) You start off by squaring both sides

TheSmartOne (thesmartone):

\(\sf\Large\sqrt{9x+67} = x + 5\) \(\sf\Large(\sqrt{9x+67})^2 = (x + 5)^2\) Simplifying that gets us: \(\sf\Large 9x+67 = (x + 5)^2\) So, expand (x+5)^2 what is it? :P

OpenStudy (anonymous):

x^2 +10x +25

TheSmartOne (thesmartone):

correct

OpenStudy (anonymous):

thats where i got stuck

OpenStudy (anonymous):

because that x^2 doesnt go anywhere

TheSmartOne (thesmartone):

\(\sf\Large 9x+67 =x^2 + 10x + 25\) now make the equation set to 0 so bring 9x + 67 on the other side :) subtract 9x and subtract 67 on both sides :)

OpenStudy (anonymous):

yeah, x^2 + x -42

OpenStudy (anonymous):

oh wait am i supposed to factor this?

jimthompson5910 (jim_thompson5910):

yes or use the quadratic formula

TheSmartOne (thesmartone):

^

OpenStudy (anonymous):

well factoring it gives me a weird number

OpenStudy (anonymous):

whats the quadratic formula?

jimthompson5910 (jim_thompson5910):

to factor, find two numbers that a) multiply to -42 (last term) AND b) add to 1 (middle coefficient)

OpenStudy (anonymous):

-6 7

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

so it factors to (x-6)(x+7)

OpenStudy (anonymous):

oh i see, i typed it wrong on my calc

OpenStudy (anonymous):

so then how is it 6?

jimthompson5910 (jim_thompson5910):

so if ` (x-6)(x+7) = 0` then either `x-6 = 0` OR `x+7 = 0`

jimthompson5910 (jim_thompson5910):

you then solve each equation for x

OpenStudy (anonymous):

if i set both to zero i will have solutions 6 and -7. but only one can be the answer i think

jimthompson5910 (jim_thompson5910):

so the *possible* solutions are x = 6 or x = -7 what you need to do now is check each possible solution back into the original equation

jimthompson5910 (jim_thompson5910):

Here's an example say we had `sqrt(x) = -1` squaring both sides gives `x = 1` but plugging that back into `sqrt(x) = -1` leads to `1 = -1` which is false. So `x = 1` is an extraneous solution (not a real solution at all)

OpenStudy (anonymous):

and -7 doesnt look like it works because it ends up being 2=-2

OpenStudy (anonymous):

so thats false and only 6 is true right?

jimthompson5910 (jim_thompson5910):

yes, only x=6 is the true solution. x=-7 is extraneous -7 satisfies every equation after you square both sides but it does NOT satisfy the original equation

OpenStudy (anonymous):

alright thanks! :)

jimthompson5910 (jim_thompson5910):

no problem